# On $L^{p}$-semigroup to Stokes equation with dynamic boundary condition in the half-space

### v2

@@ Abstract.
We consider evolutionary Stokes system, coupled with the so-called dynamic boundary condition, in the simple geometry of a $d$-dimensional half-space. Using the Fourier transform, we obtain an explicit formula for the resolvent. Maximal regularity estimates and existence of analytic semigroup in the $L^{p}$-setting are then deduced using the methods of $\mathcal{H}^{\infty}$-calculus.
@@
Stokes equations, dynamic boundary conditions, unbounded domains, maximal regularity, analytic semigroup
@@ 1. Introduction
(1) $\partial_{t}\bm{u}-\operatorname{div}2\nu D\bm{u}+\nabla\pi$ $=\bm{f},\qquad\operatorname{div}\bm{u}=0\quad\textrm{ in $(0,T)\times\Omega$},$
@@
$\beta\partial_{t}\bm{u}+\alpha\bm{u}+2\nu[{(D\bm{u})}\bm{n}]_{\tau}$ $=\beta\bm{h},\qquad\bm{u}\cdot\bm{n}=0\quad\textrm{ on $(0,T)\times\Gamma$},$
@@
where $\Omega=\{x_{d}>0\}$ is the $d$-dimensional half-space, and $\Gamma=\{x_{d}=0\}$ the corresponding boundary. Note that (1) consists of two evolution equations, coupled via the trace operator. The problem can be efficiently treated in the Hilbert $L^{2}$-setting, since there is an underlying symmetric operator $A:V\to V^{*}$, where $V$ is a suitable subspace to $W^{1,2}(\Omega)$, see e.g. [1], cf. also the weak formulation (6) below. However, for $p\neq 2$, the $L^{p}$-theory is expectedly more difficult and up to our knowledge, is so far missing in the literature. In the present paper, we intend to partially amend this, with two main results being: the existence of an analytic semigroup in $L^{p}(\Omega)\times L^{p}(\Gamma)$, and the optimal $W^{2,p}$ and $W^{1,p}$-regularity estimates for the corresponding stationary (resolvent) problem.
Although our results are formulated for the case of half-space $\Omega=\{x_{d}>0\}$ only, extension to more general domains should be possible, using standard localization and perturbation arguments. Note that this would require extension of our theorems to the case of functions $\bm{u}$ with non-zero divergence, which is rather straighforward, using the estimates of [8].
@@
The paper is organized as follows: In Section 2, we introduce function spaces and state our main results: Theorem 2.1 and Theorem 2.2, providing resolvent estimates and optimal regularity in the $L^{p}$-setting. In Corollary 2.1, we show the existence of an analytic semigroup. We also identify the domain of its generator. Main technical results of the paper then follow: in Section 3, we use (partial) Fourier transform, in the spirit of [8], to express the resolvent in terms of the boundary right-hand side. Related key estimates in $L^{p}$ spaces are then obtained in Section 4. Finally, proofs of the main theorems are completed in Section 5, essentially by combining results of previous section with the well-known $L^{p}$-theory for the (homogeneous) Dirichlet problem.
@@
Let us conclude this introduction by discussing some related, both classical and recent results. For the Stokes system subject to a homogeneous Dirichlet boundary conditions, the existence of $L^{p}$-analytic semigroup was proven in [10] via the theory of pseudo-differential operators. For a more elementary approach based on the Fourier transform and the classical multiplier theorems, see [8]. Recently, a different method using the maximal operator was proposed in [16]; however, the method seems to be limited to bounded domains.
@@
There is also an extensive literature devoted to problems related to (1), but crucially, without the time derivative in the boundary equation. For example, a fairly complete $L^{p}$-theory, including the existence of an analytic semigroup, has been established in [2], see also [9] and the references therein; yet the results again seem to be limited to bounded $\Omega$ only. In the case of a half-space, the problem was treated in [15], albeit for $\bm{h}=0$ only. A more modern approach based on the $\mathcal{H}^{\infty}$-calculus was employed here.
@@
A rather comprehensive functional setup for treating resolvent problems related to dynamic boundary conditions is developed in [7], see also the references therein. This method does provide sharp regularity estimates in fractional Sobolev spaces, yet direct application to Stokes problem – due to the incompressibility condition and the pressure term – does not seem viable.
@@ 2. Notation and main results
We set $\Omega=\{(x^{\prime},x_{d})\in\mathbb{R}^{d-1}\times\mathbb{R};x_{d}>0\}$ and identify $\Gamma=\partial\Omega$ with $\mathbb{R}^{d-1}$. Furthermore, $\bm{n}=(0,-1)$ is the outer normal, and $\tau:\mathbb{R}^{d}\to\mathbb{R}^{d-1}$ denotes the tangential component. For $p\in(1,\infty)$ and $k\geq 1$ integer, we denote by $L^{p}$ and $W^{k,p}$ the usual Lebesgue and Sobolev spaces. We also need Besov spaces $B^{s,p}_{p}$ with a general real $s$; we allow for non-integer $s$ in the spirit of [4]. The subscripts $\sigma$ or $n$ stand for the solenoidality, and zero normal component, respectively. Finally, we will extensively use the partial Fourier transform $\mathcal{F}_{x^{\prime}\to\xi}$, which is defined by
@@
(3) $\text{div}\,\bm{u}$ $=0\quad\text{ in }\Omega,$
@@
(5) $(\lambda+\alpha)\bm{u}+2[(D\bm{u})\bm{n}]_{\tau}$ $=\bm{h}\quad\text{ on }\Gamma,$
where $\lambda$ is a complex number. Note that $\operatorname{div}2D\bm{u}=\Delta\bm{u}$. Multiplying (2) by $\bm{\varphi}\in C^{\infty}_{0,\sigma}(\overline{\Omega})$ and integrating by parts, one obtains the integral formulation
(6) $\begin{gathered}\lambda\int_{\Omega}\bm{u}\cdot\bm{\varphi}\,{\rm d}x+(\lambda+\alpha)\int_{\Gamma}\bm{u}\cdot\bm{\varphi}\,{\rm d}S+\int_{\Omega}2D\bm{u}:D\bm{\varphi}\,{\rm d}x\\ =\int_{\Omega}\bm{f}\cdot\bm{\varphi}\,{\rm d}x+\int_{\Gamma}\bm{h}\cdot\bm{\varphi}\,{\rm d}S.\end{gathered}$
This problem is easy to handle in the $L^{2}$-setting, using Korn’s inequality, see [2] or [1]. However, our objective is to obtain the $L^{p}$ theory, as stated in the following two main theorems.
@@
Let $\alpha\geq 0$, $p\in(1,+\infty)$ and $\lambda$ is complex with $\mathfrak{Re}\,\lambda>0$.
@@
    Let $\bm{f}\in L^{r}_{\sigma,n}(\Omega)$, where $r$ is given by the relation $1/r=1/p+1/d$, and let $\bm{h}\in B^{-1/p,p}_{p;n}(\Gamma)$. Then one has the estimate
    (8) $\lVert\bm{u}\rVert_{W^{1,p}(\Omega)}\leq C_{1}\big{(}\lVert\bm{f}\rVert_{L^{r}(\Omega)}+\lVert\bm{h}\rVert_{B^{-1/p,p}_{p}(\Gamma)}\big{)}.$
@@
    Let $\bm{f}\in L^{p}_{\sigma,n}(\Omega)$ and $\bm{h}\in B^{1-1/p,p}_{p;n}(\Gamma)$. Then the solution to resolvent problem (2)-(5) is strong, and one has the estimate
@@ 3. Fundamental solution
Here, we deal with (2)-(5) for $\bm{f}=0$, $g=0$, i.e. we consider the system
@@
(12) $\text{div}\,\bm{u}$ $=0\quad\text{ in }\Omega,$
@@
(14) $(\lambda+\alpha)\bm{u}+2[(D\bm{u})\bm{n}]_{\tau}$ $=\bm{\phi}\quad\text{ on }\Gamma.$
@@ 3.1. Reducing the equations
By $\text{div}^{\prime}$, $\Delta^{\prime}$ we denote the corresponding operator with respect to $x^{\prime}$. Let us now simplify (13)-(14). First, $\bm{n}=(0,-1)$ and thus
@@
$2[(D\bm{u})\bm{n}]_{\tau}=2(D\bm{u})\bm{n}-2((D\bm{u})\bm{n}\cdot\bm{n})\bm{n}=-\begin{pmatrix}\partial_{d}\bm{u}^{\prime}\\ 0\end{pmatrix}.$
@@
Now, we aplly $\partial_{d}\text{div}^{\prime}$ to the first $(d-1)$ equations of (11) and $-\Delta^{\prime}$ to the last one; adding these equations together we obtain
@@
The pressure term clearly vanishes and in the first bracket we use $\text{div}\,\bm{u}=0\Leftrightarrow\text{div}^{\prime}\,\bm{u}^{\prime}=-\partial_{d}u_{d}$ to get
@@
Finally, we deal with the boundary system. We apply $-\text{div}^{\prime}$ to (16) to get
$-(\lambda+\alpha)\,\text{div}^{\prime}\,\bm{u}^{\prime}+\partial_{d}\text{div}^{\prime}\,\bm{u}^{\prime}=-\text{div}^{\prime}\,\bm{\phi},$
and thanks to $\text{div}^{\prime}\,\bm{u}^{\prime}=-\partial_{d}u_{d}$ we achieve
$(\lambda+\alpha-\partial_{d})\partial_{d}u_{d}=-\text{div}^{\prime}\,\bm{\phi}\quad\text{ on }\Gamma.$
@@
(17) $\begin{split}-(\lambda-\Delta)\Delta u_{d}&=0\quad\text{ in }\Omega,\\ (\lambda+\alpha-\partial_{d})\partial_{d}u_{d}&=-\text{div}^{\prime}\,\bm{\phi}\quad\text{ on }\Gamma,\\ u_{d}&=0\quad\text{ on }\Gamma.\end{split}$
@@ 3.2. Fourier transform
Now, we need to express $\widehat{\pi}$ and $\widehat{\bm{u}^{\prime}}$ using $m_{0}$. We start with the pressure. Let us apply $\text{div}^{\prime}$ to the first $(d-1)$ equations of (11); we get
$\Delta^{\prime}\pi=-\text{div}^{\prime}(\lambda-\Delta)\bm{u}^{\prime}\quad\text{ in }\Omega.$
Because of $\text{div}^{\prime}\,\bm{u}^{\prime}=-\partial_{d}u_{d}$ we rewrite it in the form
@@
$\text{div}\,\bm{v}^{\prime}$ $=0\quad\text{ in }\Omega,$
@@ 4. Multipliers and $L^{p}$-bounds
see e.g. [11, Theorem 5.2.7]. It follows once we prove
@@ 5. Proof of the main results
$\text{div}\,\bm{u}^{1}$ $=0\quad\text{ in }\Omega,$
@@
$\text{div}\,\bm{u}^{2}$ $=0\quad\text{ in }\Omega,$
@@
$(\lambda+\alpha)\bm{u}^{2}+2[(D\bm{u}^{2})\bm{n}]_{\tau}$ $=\bm{\phi}\quad\text{ on }\Gamma,$
@@
(41) $\bm{\phi}=\bm{h}-2[(D\bm{u}^{1})\bm{n}]_{\tau}.$
Assume that $\lambda\in\mathbb{C}$ with $|\lambda|>\omega$, where $\omega>0$ is fixed. Due to the standard theory (see e.g. [8, Theorem 1.2]) we know that
@@
$||\bm{\phi}||_{L^{p}(\Gamma)}$ $=||\bm{h}-2[(D\bm{u}^{1})\bm{n}]_{\tau}||_{L^{p}(\Gamma)}\leq c||(\bm{f},\bm{h})||_{{L^{p}(\Omega)}\times{L^{p}(\Gamma)}}.$
@@
We start with part (ii). Employing again the splitting (3941), regularity of $\bm{u}^{1}$ is proven in [10] or [8]. Note that $\bm{\phi}\in B^{1-1/p,p}_{p}(\Gamma)$ by the trace theorem [3, Theorem 7.39].
@@
$\widehat{\pi}(\xi,y)=f(z)e^{-\eta zy}\widehat{\nabla^{\prime}\bm{v}}(\xi,y)$
@@
It remains to show part (i). We note first that $\bm{u}^{1}\in W^{2,r}(\Omega)$, and thus $\nabla\bm{u}^{1}\in W^{1,p}(\Omega)$. We also have $\nabla\pi^{1}\in L^{r}(\Omega)$. The key step is to bound $D\bm{u}^{1}\in B^{-1/p,p}_{p}(\Gamma)$, to which we need the following auxiliary result.
@@
Let $\bm{U}\in L^{p}(\Omega)$ and $\operatorname{div}\bm{U}\in L^{r}(\Omega)$, where $1/r=1/p+1/d$. Then $(\bm{U}\cdot\bm{n})_{\tau}\in B^{-1/p,p}_{p;n}(\Gamma)$.
@@
and note that the right-hand side can be bounded in terms of $\bm{\varphi}\in W^{1,p^{\prime}}_{n}(\Omega)$. It remains to observe that
$\gamma(W^{1,p^{\prime}}_{n}(\Omega))=B^{1/p,p^{\prime}}_{p^{\prime};n}=(B^{-1/p,p}_{p;n}(\Gamma))^{*}\,.$

### v3

# On $L^{p}$-semigroup to Stokes equation with dynamic slip boundary condition in the half-space

@@ Abstract.
We consider evolutionary Stokes system, coupled with the so-called dynamic slip boundary condition, in the simple geometry of a $d$-dimensional half-space. Using the Fourier transform, we obtain an explicit formula for the resolvent. Optimal regularity estimates and existence of analytic semigroup in the $L^{p}$-setting are then deduced using the methods of $\mathcal{H}^{\infty}$-calculus.
@@
Stokes equations, dynamic slip boundary condition, unbounded domains, optimal regularity, analytic semigroup
@@ 1. Introduction
(1) $\partial_{t}\bm{u}-\operatorname{div}2\nu{\rm D}\bm{u}+\nabla\pi$ $=\bm{f},\qquad\operatorname{div}\bm{u}=0\quad\textrm{ in $(0,T)\times\Omega$},$
@@
$\beta\partial_{t}\bm{u}+\alpha\bm{u}+2\nu[{({\rm D}\bm{u})}\bm{n}]_{\tau}$ $=\beta\bm{h},\qquad\bm{u}\cdot\bm{n}=0\quad\textrm{ on $(0,T)\times\Gamma$},$
@@
where $\Omega=\{x_{d}>0\}$ is the $d$-dimensional half-space, and $\Gamma=\{x_{d}=0\}$ the corresponding boundary. Note that (1) consists of two evolution equations, coupled via the trace operator. The problem can be efficiently treated in the Hilbert $L^{2}$-setting, since there is an underlying symmetric operator $A:V\to V^{*}$, where $V$ is a suitable subspace to $W^{1,2}(\Omega)$, see e.g. [1], cf. also the weak formulation (6) below. However, the $L^{p}$-theory is expectedly more difficult if $p\neq 2$ and up to our knowledge, the problem (1) has not been treated in the literature (although a number of closely related results are available, see also the discussion below). In the present paper, we intend to contribute to this, with two main results being: the existence of an analytic semigroup in $L^{p}(\Omega)\times L^{p}(\Gamma)$, and the optimal $W^{2,p}$ and $W^{1,p}$-regularity estimates for the corresponding stationary (resolvent) problem.
@@
Our paper is organized as follows. In Section 2, we introduce function spaces and state two main results: Theorem 2.1 and Theorem 2.2, providing resolvent estimates of optimal regularity both for weak and strong solutions. In Corollary 2.1, we show the existence of an analytic semigroup; we also identify the domain of its generator. Main technical results of the paper follow: in Section 3, we use (partial) Fourier transform, in the spirit of [9], to express the resolvent in terms of the boundary right-hand side. Key estimates in $L^{p}$ spaces are then obtained in Section 4. Our approach remains more elementary in that we only rely on the classical Mikhlin multiplier theorem. One can however note that Lemma 4.1 in fact asserts the existence of $\mathcal{H}^{\infty}$-calculus for the operator $(-\Delta)^{1/2}$. Finally, proofs of the main theorems are completed in Section 5, essentially by combining results of Section 4 with the well-known $L^{p}$-theory for the (homogeneous) Dirichlet problem, as given e.g. in [10].
Although our results are formulated for the case of half-space $\Omega=\{x_{d}>0\}$ only, extension to more general domains is straighforward, using standard localization and perturbation arguments. Crucially, this requires extension of the estimates to the case of functions $\bm{u}$ with non-zero divergence. We treat this situation explicitly in part (iii) of Theorem 2.2.
@@
Let us conclude the introduction by discussing some related, both classical and more recent, results. Concerning the Stokes system subject to a homogeneous Dirichlet boundary condition, the existence of $L^{p}$-analytic semigroup was proven in [12] via the theory of pseudo-differential operators. For a more elementary approach based on the Fourier transform and classical multiplier theorems, see [9]. Recently, a different method using the maximal operator was proposed in [18]; however, the method seems to be limited to bounded domains.
@@
There is also an extensive literature devoted to problems related to (1), but crucially, without the time derivative in the boundary equation. For example, a fairly complete $L^{p}$-theory, including the existence of an analytic semigroup, has been established in [2], see also [11] and the references therein; yet the results again seem to be limited to bounded $\Omega$ only. In the case of a half-space, the problem was treated in [17], albeit for $\bm{h}=0$ only. A more modern approach based on the $\mathcal{H}^{\infty}$-calculus was employed here. More recently, a closely related class of problems with dynamic boundary conditions was treated in [6], and maximal regularity results were obtained.
@@
A rather comprehensive functional setup for treating resolvent problems related to dynamic boundary conditions is described in [8]; see also the references therein. This method does provide sharp regularity estimates in fractional Sobolev spaces, yet direct application to Stokes problem – due to the incompressibility condition and the pressure term – does not seem viable.
@@ 2. Notation and main results
We set $\Omega=\{(x^{\prime},x_{d})\in\mathbb{R}^{d-1}\times\mathbb{R};x_{d}>0\}$ and identify $\Gamma=\partial\Omega$ with $\mathbb{R}^{d-1}$. Furthermore, $\bm{n}=(0,-1)$ is the outer normal, and $\tau:\mathbb{R}^{d}\to\mathbb{R}^{d-1}$ denotes the tangential component. For $p\in(1,\infty)$ and $k\geq 1$ integer, we denote by $L^{p}$ and $W^{k,p}$ the usual Lebesgue and Sobolev spaces. Negative Sobolev space $W^{-1,p}$ is defined as the completion of $L^{q}$ with respect to the norm
$\lVert\bm{u}\rVert_{W^{-1,p}(\Omega)}=\sup\Big{\{}\int_{\Omega}\bm{u}\cdot\bm{\varphi}\,{\rm d}x;\ \lVert\bm{\varphi}\rVert_{W^{1,p^{\prime}}(\Omega)}\leq 1\Big{\}}.$
It it well-known (see e.g. [3, Section 3.14]) that $W^{-1,p}$ can be identified with the dual to $W^{1,p^{\prime}}$.
We also need Besov spaces $B^{s,p}_{p}$ with a general real $s$; we allow for non-integer $s$ in the spirit of [4]. The subscripts $\sigma$ or $n$ stand for the solenoidality, and zero normal component, respectively. Finally, we will extensively use the partial Fourier transform $\mathcal{F}_{x^{\prime}\to\xi}$, which is defined by
@@
(3) $\operatorname{div}\bm{u}$ $=0\quad\text{ in }\Omega,$
@@
(5) $(\lambda+\alpha)\bm{u}+2[({\rm D}\bm{u})\bm{n}]_{\tau}$ $=\bm{h}\quad\text{ on }\Gamma,$
where $\lambda$ is a complex number. Note that $\operatorname{div}2{\rm D}\bm{u}=\Delta\bm{u}$. Multiplying (2) by $\bm{\varphi}\in C^{\infty}_{0,\sigma}(\overline{\Omega})$ and integrating by parts, one obtains the integral formulation
(6) $\begin{gathered}\lambda\int_{\Omega}\bm{u}\cdot\bm{\varphi}\,{\rm d}x+(\lambda+\alpha)\int_{\Gamma}\bm{u}\cdot\bm{\varphi}\,{\rm d}S+\int_{\Omega}2{\rm D}\bm{u}:{\rm D}\bm{\varphi}\,{\rm d}x\\ =\int_{\Omega}\bm{f}\cdot\bm{\varphi}\,{\rm d}x+\int_{\Gamma}\bm{h}\cdot\bm{\varphi}\,{\rm d}S.\end{gathered}$
This problem is easy to handle in the $L^{2}$-setting, using Korn’s inequality, see [2] or [1]. However, our objective is to obtain the $L^{p}$-theory, as stated in the following two main theorems.
@@
Let $\alpha\geq 0$, $p\in(1,+\infty)$ and $\lambda\in\mathbb{C}$ with $\mathfrak{Re}\,\lambda>0$.
@@
    Let $\bm{f}\in W^{-1,p}(\Omega)$, and let $\bm{h}\in B^{-1/p,p}_{p;n}(\Gamma)$. Then one has the estimate
    (8) $\lVert\bm{u}\rVert_{W^{1,p}(\Omega)}\leq C_{1}\big{(}\lVert\bm{f}\rVert_{W^{-1,p}(\Omega)}+\lVert\bm{h}\rVert_{B^{-1/p,p}_{p}(\Gamma)}\big{)}.$
@@
    Let $\bm{f}\in L^{p}_{\sigma,n}(\Omega)$ and $\bm{h}\in B^{1-1/p,p}_{p;n}(\Gamma)$. Then the solution to resolvent problem (2)–(5) is strong, and one has the estimate
@@
  • (iii)
    If (3) is generalized to $\operatorname{div}\bm{u}=g$, the estimates remain valid, with the norm of $g$ in $L^{p}(\Omega)$ or $W^{1,p}(\Omega)$ added to the right-hand side of (8) or (9), respectively.
@@ 3. Fundamental solution
Here, we deal with (2)–(5) for $\bm{f}=0$, i.e. we consider the system
@@
(12) $\operatorname{div}\bm{u}$ $=0\quad\text{ in }\Omega,$
@@
(14) $(\lambda+\alpha)\bm{u}+2[({\rm D}\bm{u})\bm{n}]_{\tau}$ $=\bm{\phi}\quad\text{ on }\Gamma.$
@@ 3.1. Reducing the equations
By $\operatorname{div}^{\prime}$, $\Delta^{\prime}$ we denote the corresponding operator with respect to $x^{\prime}$. Let us now simplify (13)–(14). First, $\bm{n}=(0,-1)$ and thus
@@
$2[({\rm D}\bm{u})\bm{n}]_{\tau}=2({\rm D}\bm{u})\bm{n}-2(({\rm D}\bm{u})\bm{n}\cdot\bm{n})\bm{n}=-\begin{pmatrix}\partial_{d}\bm{u}^{\prime}\\ 0\end{pmatrix}.$
@@
Now, we aplly $\partial_{d}\operatorname{div}^{\prime}$ to the first $(d-1)$ equations of (11) and $-\Delta^{\prime}$ to the last one; adding these equations together we obtain
@@
The pressure term clearly vanishes and in the first bracket we use $\operatorname{div}\bm{u}=0\Leftrightarrow\\ diver^{\prime}\bm{u}^{\prime}=-\partial_{d}u_{d}$ to get
@@
Finally, we deal with the boundary system. We apply $-\operatorname{div}^{\prime}$ to (16) to get
$-(\lambda+\alpha)\,\operatorname{div}^{\prime}\bm{u}^{\prime}+\partial_{d}\operatorname{div}^{\prime}\bm{u}^{\prime}=-\operatorname{div}^{\prime}\bm{\phi},$
and thanks to $\operatorname{div}^{\prime}\bm{u}^{\prime}=-\partial_{d}u_{d}$ we achieve
$(\lambda+\alpha-\partial_{d})\partial_{d}u_{d}=-\operatorname{div}^{\prime}\bm{\phi}\quad\text{ on }\Gamma.$
@@
(17) $\begin{split}-(\lambda-\Delta)\Delta u_{d}&=0\quad\text{ in }\Omega,\\ (\lambda+\alpha-\partial_{d})\partial_{d}u_{d}&=-\operatorname{div}^{\prime}\bm{\phi}\quad\text{ on }\Gamma,\\ u_{d}&=0\quad\text{ on }\Gamma.\end{split}$
@@ 3.2. Fourier transform
Now, we need to express $\widehat{\pi}$ and $\widehat{\bm{u}^{\prime}}$ using $m_{0}$. We start with the pressure. Let us apply $\operatorname{div}^{\prime}$ to the first $(d-1)$ equations of (11); we get
$\Delta^{\prime}\pi=-\operatorname{div}^{\prime}(\lambda-\Delta)\bm{u}^{\prime}\quad\text{ in }\Omega.$
Because of $\operatorname{div}^{\prime}\bm{u}^{\prime}=-\partial_{d}u_{d}$ we rewrite it in the form
@@
$\operatorname{div}\bm{v}^{\prime}$ $=0\quad\text{ in }\Omega,$
@@ 4. Multipliers and $L^{p}$-bounds
see e.g. [13, Theorem 5.2.7]. It follows once we prove
@@ 5. Proof of the main results
$\operatorname{div}\bm{u}^{1}$ $=0\quad\text{ in }\Omega,$
@@
$\operatorname{div}\bm{u}^{2}$ $=0\quad\text{ in }\Omega,$
@@
$(\lambda+\alpha)\bm{u}^{2}+2[({\rm D}\bm{u}^{2})\bm{n}]_{\tau}$ $=\bm{\phi}\quad\text{ on }\Gamma,$
@@
(41) $\bm{\phi}=\bm{h}-2[({\rm D}\bm{u}^{1})\bm{n}]_{\tau}.$
Assume that $\lambda\in\mathbb{C}$ with $|\lambda|>\omega$, where $\omega>0$ is fixed. Due to the standard theory (see e.g. [9, Theorem 1.2]) we know that
@@
$||\bm{\phi}||_{L^{p}(\Gamma)}$ $=||\bm{h}-2[({\rm D}\bm{u}^{1})\bm{n}]_{\tau}||_{L^{p}(\Gamma)}\leq c||(\bm{f},\bm{h})||_{{L^{p}(\Omega)}\times{L^{p}(\Gamma)}}.$
@@
We split the solution $\bm{u}=\bm{u}^{1}+\bm{u}^{2}$ as above, taking moreover $\operatorname{div}\bm{u}^{1}=g$ in (39). Hence, part (iii) of the Theorem is taken into account.
Let us now consider the strong estimates, i.e. part (ii). Regularity of $\bm{u}^{2}$ follows as in [9], see also [10, Theorem IV.3.2]. Note that $\bm{\phi}\in B^{1-1/p,p}_{p}(\Gamma)$ by the trace theorem [3, Theorem 7.39].
@@
$\widehat{\pi}(\xi,y)=f(z)e^{-\eta zy}\widehat{\nabla^{\prime}\bm{v}}(\xi,y).$
@@
It remains to show part (i). Existence and regularity of $\bm{u}^{1}$ follows from [10, Theorem IV.3.3]. The key step is to bound ${\rm D}\bm{u}^{1}\in B^{-1/p,p}_{p}(\Gamma)$, to which we need the following auxiliary result.
@@
Let $\bm{U}\in L^{p}(\Omega)$ and $\operatorname{div}\bm{U}\in W^{-1,p}(\Omega)$, Then $(\bm{U}\cdot\bm{n})_{\tau}\in B^{-1/p,p}_{p;n}(\Gamma)$.
@@
and note that the right-hand side can be bounded in terms of $\bm{\varphi}\in W^{1,p^{\prime}}_{n}(\Omega)$. Denoting formally by $\gamma$ the trace operator, it remains to observe that
$\gamma(W^{1,p^{\prime}}_{n}(\Omega))=B^{1/p,p^{\prime}}_{p^{\prime};n}(\Gamma)=(B^{-1/p,p}_{p;n}(\Gamma))^{*}\,.$